# Why Can’t I Open My Drawer? Mitigating Object-Driven Shortcuts in Zero-Shot Compositional Action Recognition

### v1

@@ 3.3 Empirical Diagnosis of Object-driven Shortcuts in ZS-CAR
In this section, we systematically diagnose the key failure mode in ZS-CAR—object-driven shortcuts. We demonstrate that these shortcuts prevent ZS-CAR models from learning robust verb representations, consequently leading to poor generalization to unseen compositions. We present additional evidence of object-driven shortcut learning in ZS-CAR models in Section 5 and the supplementary material.
Objects are easier to learn than verbs.
To investigate the learning difficulty of verbs and objects, we train a randomly initialized model on a subset of Sth-com [16]. Based on the observation that the easier-to-learn cue is captured earlier and exhibits higher accuracy [25], we can analyze the learning difficulty of each component. In Figure 2 (a), the results show a clear trend: the model learns objects faster and consistently achieves higher accuracy for objects than for verbs. While intuitively expected, these results empirically verify that objects are inherently easier to learn than verbs in compositional action recognition.
Object-driven shortcuts in ZS-CAR
Under strong co-occurrence priors, models choose shortcuts emerged from easier-to-learn cues, rather than learn robust representations [25, 30, 11, 3, 17, 6]. We argue this shortcut learning behavior accelerates in ZS-CAR due to the asymmetric learning difficulty. To verify this, inspired by prior work [30], we train models on a perfectly biased training set and test on two distinct splits—bias-aligned seen compositions and bias-conflict unseen compositions— utilizing Sth-com [16]. As shown in Figure 2 (b), in the both cases of the randomly initialized ViT [10] and CLIP [28], the object accuracies for unseen compositions increased as the seen composition accuracies also increased, while the verb accuracies for unseen compositions drop under the random accuracy. Paradoxically, CLIP [16], well-known for its superior representational power, exhibits the highest object accuracy (94.0%) and the lowest verb accuracy (1.9%) concurrently on unseen compositions. These results demonstrate models tend to fall into object-driven shortcuts to exploit co-occurrence priors in the training data, resulting in poor verb representations.
Object-driven shortcuts directly harm unseen generalization.
Next, we investigate whether the state-of-the-art (SOTA) model also fails to generalize to unseen compositions as a consequence of these object-driven shortcuts. As shown in Figure 3, we trace the validation accuracies along with FSP and FCP ratios—metrics designed to quantify the model’s reliance on the training statistics—of C2C [16] throughout the training progress on Sth-com [16]. The clear correlation between increased FSP/FCP ratios and widen seen-unseen accuracy gap indicates that the model overfits to the training statistics.
Table 1: The SOTA model exhibits a negative Compositional Gap on unseen compositions. We measure Compositional Gap ($\Delta_{\mathrm{CG}}$) of the SOTA ZS-CAR model. We employ C2C [16] trained on Sth-com [16] with 30 epochs and evaluate the model in open-world setting on the Sth-com [16] test set. The model shows a negative $\Delta_{\mathrm{CG}}$ on unseen compositions, indicating its composition accuracy falls below the level expected from independent verb–object predictions.
Split $\mathrm{Acc}^{V}$ $\mathrm{Acc}^{O}$ $\mathrm{Acc}^{V}\times\mathrm{Acc}^{O}$ $\mathrm{Acc}^{C}$ $\Delta_{\text{CG}}$
Seen 63.60 67.72 43.07 46.31 +3.24
Unseen 54.36 56.10 30.50 30.08 -0.42
Compositional Gap reveals the SOTA model fails to benefit from compositional modeling.
We evaluate the Compositional Gap ($\Delta_{\text{CG}}$) performance of current SOTA ZS-CAR model [16]. As shown in Table 1, the model shows a negative $\Delta_{\text{CG}}$ on unseen compositions in Sth-com [16]. This demonstrates that, without addressing these shortcuts, the SOTA model gains no advantage from compositional modeling compared to independent modeling for unseen compositions.
@@ 4 RCORE
We introduce RCORE, a learning framework that mitigates object-driven shortcut learning in zero-shot compositional action recognition. RCORE improves compositional generalization by strengthening verb–object representations under sparse supervision through two complementary components: (i) VOCAMix, a composition-aware augmentation that expands the variety of verb–object combinations, and (ii) TORC, a temporal order regularization loss that promotes temporally grounded verb learning over static object cues. Following prior work [16], we use AIM [37], a CLIP-based [28] parameter-efficient spatio-temporal encoder, as our backbone. Figure 4 (a) gives an overview of RCORE.
Feature extraction.
@@
Given an augmented video $\hat{\mathbf{X}}$ with $T$ frames, the backbone encoder outputs frame-level features $\mathbf{f}_{t}\in\mathbb{R}^{D}$. We then transform them into verb features $\mathbf{F}^{V}\in\mathbb{R}^{T\times D}$ and an object features $\mathbf{f}^{O}\in\mathbb{R}^{D}$ using dedicated verb and object encoders. We obtain text embeddings for verbs and objects, $\mathbf{e}^{V}$ and $\mathbf{e}^{O}$, by feeding class-specific prompts (e.g., “a photo of {verb/object}) into the text encoder.
4.1 Composition-Aware Augmentation
To alleviate the sparsity of compositional training data, we propose Verb–Object Composition-Aware Augmentation (VOCAMix), which synthesizes plausible but unseen combinations while preserving the temporal structure of the primary video.
As shown in Figure 4 (b), VOCAMix replaces the motion region of a video with a static object region from another sample in the mini-batch. Given a primary video $\mathbf{X}_{i}$ with label $(\mathbf{y}_{i}^{V},\mathbf{y}_{i}^{O})$ and a randomly selected video $\mathbf{X}_{j}$ with object label $\mathbf{y}_{j}^{O}$ ($i\neq j$), we construct an augmented video $\hat{\mathbf{X}}_{i}$ and a soft object label:
@@ 4.1 Composition-Aware Augmentation
$h(\hat{\mathbf{X}}_{i}(k,:))$ $=(1-\lambda)\,h(\mathbf{X}_{i}(k,:))+\lambda\,h(\mathbf{X}_{j}(\lfloor T/2\rfloor,:)),$ (2)
$\tilde{\mathbf{y}}^{O}_{i}$ $=(1-\lambda)\,\mathbf{y}_{i}^{O}+\lambda\,\mathbf{y}_{j}^{O},$ (3)
where $h(\cdot)$ is a high motion-region estimation function, $k\in\{1,\ldots,T\}$ is the frame index, and $\lambda\in[0,1]$ controls the strength of object injection. We adopt a simple, learning-free estimator [9] for $h(\cdot)$. We keep the verb label $\mathbf{y}_{i}^{V}$ unchanged to preserve temporal semantics. By injecting a consistent static object across the temporal axis, VOCAMix preventing disruption of temporal cues [33] while expanding compositional diversity. while synthesizing novel compositions.
4.2 Temporal Order Regularization Loss
To counter object-driven shortcuts in verb learning, we introduce the Temporal Order Regularization for Composition (TORC) loss. TORC loss forces the model to rely on temporal structure: it penalizes alignment with temporally incorrect feature sequences and suppresses confident verb predictions when temporal ordering is corrupted. This regularization steers the model away from static object cues and strengthens temporally grounded verb representations.
Figure 5: Analysis on the effects of RCORE on the Sth-com [16] dataset. (a) RCORE prevents the False Co-occurrence Prediction (FCP) ratio from increasing during training, whereas the baseline shows a clear rise in FCP. As a result, RCORE consistently maintains a smaller seen–unseen accuracy gap ($\Delta_{SU}$) throughout training. (b) The cosine similarity between the original and reversed verb features becomes strongly negative for RCORE as training progresses, indicating improved temporal discriminative capability. In contrast, the baseline maintains a high similarity (0.91), revealing limited temporal sensitivity. (c) On the Temporal subset, RCORE exhibits a substantially larger performance gap between original and temporally shuffled features compared to the baseline, demonstrating that RCORE learns verb representations that depend on temporal dynamics rather than static object cues. Best viewed with zoom and color.
Temporal perturbation.
@@ 4.2 Temporal Order Regularization Loss
Given frame-level features $\mathbf{F}=(\mathbf{f}_{1},\dots,\mathbf{f}_{T})$, we construct two perturbed variants: (i) a reversed sequence $\mathbf{F}_{\text{rev}}=(\mathbf{f}_{T},\dots,\mathbf{f}_{1})$, and (ii) a temporally shuffled sequence $\mathbf{F}_{\text{shuffled}}=\pi(\mathbf{F})$, where $\pi$ is a random permutation sampled from $\mathcal{P}$. We then feed $\mathbf{F}_{\text{rev}}$ and $\mathbf{F}_{\text{shuffled}}$ into the verb encoder to obtain perturbed verb features $\mathbf{f}^{V}_{\text{rev}}\in\mathbb{R}^{D}$ and $\mathbf{f}^{V}_{\text{shuffled}}\in\mathbb{R}^{D}$.
TORC loss.
We first push the model to distinguish forward and reversed temporal semantics. We minimize the cosine similarity between the original verb feature $\mathbf{f}^{V}$ and the reversed feature $\mathbf{f}^{V}_{\text{rev}}$:
$L_{\cos}=\cos(\mathbf{f}^{V},\mathbf{f}^{V}_{\text{rev}}),\vskip-5.0pt$ (4)
@@
where $\cos(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}^{\top}\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$. Temporal reversal often flips action meaning (e.g., opening vs. closing), yet the state-of-the-art model [16] yields a high cosine similarity ($0.91$; see Figure 5), indicating weak temporal discrimination.
@@
Next, we regularize the model to avoid confident verb predictions when temporal structure is disrupted. Given the set of verb text embeddings $\mathbb{E}^{V}=\{\mathbf{e}^{V}_{i}\}_{i=1}^{|\mathbb{Y}^{V}|}\in\mathbb{R}^{|\mathbb{Y}^{V}|\times D}$, we define the negative entropy loss $L_{\text{ent}}$ as:
$L_{\text{ent}}$ $=\sum_{i=1}^{|\mathbb{Y}^{V}|}p_{i}\log p_{i},\vskip-5.0pt$ (5)
@@
where $p_{i}$ denotes the predicted probability of the $i$-th verb class based on the cosine similarity between visual and textual features, and $\tau$ is a temperature parameter. Maximizing entropy prevents the model from inferring verbs from static spatial cues alone, enforcing reliance on temporal evidence. We define the TORC loss as: $L_{\text{TORC}}=L_{\cos}+L_{\text{ent}}$.
Component loss.

### v2

@@ 3.3 Empirical Diagnosis of Object-driven Shortcuts in ZS-CAR
In this section, we empirically diagnose object-driven shortcut behavior in ZS-CAR and analyze how it correlates with weak verb learning and poor generalization to unseen compositions. We show that this behavior persists even with a modern video-pretrained VLM backbone [wang2024iv2], and provide additional evidence in Section 5 and Section 0.D.
Objects are easier to learn than verbs. To compare the learning difficulty of verbs and objects, we train a randomly initialized ViT [dosovitskiy202vit] on a controlled subset of Sth-com [li2024c2c]. Specifically, we construct a balanced $10\times 10$ subset containing all 100 verb–object pairs to isolate learning dynamics from compositional sparsity. Based on prior observations that models tend to fit easier cues earlier under limited supervision [nam2020lff, scimeca2022shortcut], we compare the learning curves of verb vs. object prediction. In Figure 2 (a), the model learns objects faster and reaches higher accuracy than verbs under this controlled setting, suggesting that object prediction is easier than verb prediction in our ZS-CAR setup.
Object-driven shortcuts indeed exist in ZS-CAR. Under skewed co-occurrence statistics, models may rely on easier cues as shortcuts instead of learning the intended evidence [nam2020lff, scimeca2022shortcut, geirhos2018cuecon, bahng2019rebias, li2018resound, whycantchoi]. We test whether CLIP exhibits such object-driven shortcuts by constructing a bias-controlled training split where each verb is strongly correlated with a particular object (see construction details in Section 0.C), and evaluating on two splits: (i) bias-aligned seen compositions and (ii) bias-conflict unseen compositions. As shown in Figure 2 (b), CLIP achieves high object accuracy on unseen compositions while verb accuracy on unseen compositions drops below the uniform-chance level ($1/|\mathbb{Y}^{V}|$), indicating that object cues dominate verb prediction under this biased supervision. We observe this trend even with random initialization, implying shortcuts stem from the training distribution and objective, not just pretraining. See Section 0.C for more details.
Shortcut reliance correlates with poor unseen generalization. Next, we test whether shortcut reliance correlates with weak unseen generalization in a standard SOTA pipeline. Figure 3 plots validation accuracies together with FSP/FCP for C2C [li2024c2c] on Sth-com [li2024c2c]. As shown in Figure 3 (a), with a CLIP backbone [radford2021clip], increasing FSP/FCP accompanies a widening seen–unseen gap, consistent with growing reliance on co-occurrence statistics. Our error breakdown in Figure 3 (c) further shows that, among unseen samples that collapse to seen pairs, a large fraction corresponds to verb-collapse, supporting the presence of object-driven shortcut behavior. As shown in Figure 3 (b) and (c), with an InternVideo2 [wang2024iv2] backbone, the trends are more stable; however, FSP remains high, and the FCP breakdown still exhibits substantial verb-collapse, suggesting that stronger video pre-training alone is insufficient to eliminate the tendency to overfit to seen pairs via co-occurrence-driven shortcuts.
Table 1: $\Delta_{\mathrm{CG}}{}$ on Sth-com.
Backbone Split $\Delta_{\text{CG}}$
CLIP Seen +3.24
Unseen -0.42
InternVideo2 Seen +2.24
Unseen -0.63
$\Delta_{\text{CG}}$ indicates degraded compositional behavior on unseen compositions. We report $\Delta_{\text{CG}}$ of C2C [li2024c2c] with CLIP [radford2021clip] and InternVideo2 [wang2024iv2] backbones. As shown in Table 1, the model exhibits negative $\Delta_{\text{CG}}$ on unseen compositions even with the video-pretrained backbone, indicating that $\mathrm{Acc}^{C}$ falls below the independent reference $\mathrm{Acc}^{V}\times\mathrm{Acc}^{O}$. Together with the training-bias metrics (FSP/FCP), these results support that shortcut reliance is closely associated with weak unseen compositional generalization in current ZS-CAR pipelines.
@@ 4 RCORE
We introduce RCORE, a diagnosis-driven learning framework that mitigates object-driven shortcuts in ZS-CAR. RCORE improves compositional generalization by strengthening verb–object representations under sparse supervision through two complementary components: (i) Co-occurrence Prior Regularization (CPR) expands supervision over novel compositions and leverages frequently seen pairs as hard negatives to suppress co-occurrence priors, and (ii) Temporal Order Regularization for Composition (TORC) loss promotes temporally grounded verb learning over static object cues. Following prior work [li2024c2c, jung2025crr], we employ an adapter-based tuning, AIM [yang2023aim] for CLIP [radford2021clip] and LoRA [hu2022lora] for InternVideo2 [wang2024iv2], as our backbone. Figure 4 (a) gives an overview of RCORE.
@@
Feature extraction. Given a video $\mathbf{X}$ with $T$ frames, the backbone encoder outputs frame-level features $\mathbf{F}\in\mathbb{R}^{T\times D}$. We transform them into verb features $\mathbf{F}^{V}\in\mathbb{R}^{T\times D}$ and object features $\mathbf{f}^{O}\in\mathbb{R}^{D}$ using dedicated encoders. We obtain text embeddings for verbs and objects, $\mathbf{E}^{V}\in\mathbb{R}^{|\mathbb{Y}^{V}|\times D}$ and $\mathbf{E}^{O}\in\mathbb{R}^{|\mathbb{Y}^{O}|\times D}$, by feeding class-specific prompts into the text encoder.
4.1 CPR: Co-occurrence Prior Regularization
To mitigate object-driven shortcuts under sparse and skewed compositional supervision, we propose CPR, a training strategy that injects supervision for synthesized verb–object pairs while preserving the stability of closed-world optimization. CPR combines (i) synthesized composition supervision, (ii) a margin-based regularizer that down-weights frequently seen hard negatives, and (iii) a batch-adaptive expansion of the composition label space to make these new pairs trainable without optimizing over the full $\mathbb{Y}^{V}\times\mathbb{Y}^{O}$ space.
Figure 4: Overview of RCORE. (a) Overview of our proposed RCORE framework. (b) CPR synthesizes plausible yet unseen verb–object compositions and penalizes frequently seen hard negatives by a margin-based regularizer. (c) TORC penalizes alignment between original and temporally perturbed feature vectors, enforcing explicit temporal order modeling and reducing object-driven shortcuts.
Synthesized composition supervision. To provide explicit supervision for compositions absent from the training set, we first synthesize videos that represent new composition labels. Given a training sample $\mathbf{X}_{i}$ with composition label $\mathbf{y}_{i}^{C}=(\mathbf{y}_{i}^{V},\mathbf{y}_{i}^{O})$, we construct a new video $\tilde{\mathbf{X}}_{i}$ by injecting the static object cue from another video $\mathbf{X}_{j}$ (with $\mathbf{y}_{j}^{O}\neq\mathbf{y}_{i}^{O}$) into the high-motion regions of $\mathbf{X}_{i}$, as shown in Figure 4 (b). Specifically, for each frame $k\in\{1,\ldots,T\}$ we obtain
@@ 4.1 CPR: Co-occurrence Prior Regularization
$\vskip-3.00003pt\tilde{\mathbf{X}}_{i}^{k}$ $=\bigl(\mathbf{1}-\lambda\,\mathbf{M}_{i}^{k}\bigr)\odot\mathbf{X}_{i}^{k}\;+\;\bigl(\lambda\,\mathbf{M}_{i}^{k}\bigr)\odot\mathbf{X}_{j}^{\lfloor T/2\rfloor},\vskip-3.00003pt$ (3)
where $\lambda\in[0,1]$ controls the injection strength and $\mathbf{M}_{i}^{(k)}\in\{0,1\}^{H\times W}$ is the high-motion region mask for frame $k$, extracted by a learning-free estimator [ding2022fame]. The operator $\odot$ denotes element-wise multiplication, with $\mathbf{M}_{i}^{k}$ broadcast across channels. Accordingly, we use soft labels for the synthesized video as follows:
$\vskip-3.00003pt\tilde{\mathbf{y}}^{O}_{i}$ $=(1-\lambda)\,\mathbf{y}_{i}^{O}+\lambda\,\mathbf{y}_{j}^{O},$ (4)
$\tilde{\mathbf{y}}^{C}_{i}$ $=(1-\lambda)\,\phi(\mathbf{y}_{i}^{V},\mathbf{y}_{i}^{O})+\lambda\,\phi(\mathbf{y}_{i}^{V},\mathbf{y}_{j}^{O}),\vskip-3.00003pt$ (5)
where $\phi:\mathbb{Y}^{V}\times\mathbb{Y}^{O}\rightarrow\{0,1\}^{|\mathbb{Y}^{V}\times\mathbb{Y}^{O}|}$ maps a composition $(\mathbf{y}^{V},\mathbf{y}^{O})$ to its one-hot vector in the flattened joint label space.
Co-occurrence prior regularization loss. Even after adding supervision for synthesized compositions, skewed co-occurrence statistics can still make the model over-score a small set of frequently seen pairs. To counter this score collapse, we enforce a margin constraint that makes the target synthesized composition $\mathbf{y}^{C}_{i,\text{new}}=(\mathbf{y}_{i}^{V},\mathbf{y}_{j}^{O})$ outrank frequent seen hard negatives by at least $m$. We define the hard-negative label set for $\mathbf{y}^{C}_{i,\text{new}}$ as follows:
$\vskip-5.0pt\mathbb{Y}_{\text{hard},i}=\left\{(\mathbf{y}^{V},\mathbf{y}^{O})~\middle|~\big[(\mathbf{y}^{V}=\mathbf{y}_{i}^{V})\oplus(\mathbf{y}^{O}=\mathbf{y}_{j}^{O})\big]\setminus(\mathbf{y}^{V},\mathbf{y}^{O})\neq(\mathbf{y}_{i}^{V},\mathbf{y}_{i}^{O})\right\},\vskip-5.0pt$ (6)
where $\oplus$ denotes the exclusive or operator. Let $\mathbb{Y}_{\text{freq}}\subset\mathbb{Y}_{\text{seen}}$ be the set of frequently seen compositions (details about the criterion in Section 0.C.2). Then we define the CPR loss with margin $m$ over the penalty set $\mathbb{Y}_{\text{pen}}=\mathbb{Y}_{\text{hard},i}\cap\mathbb{Y}_{\text{freq}}$ as:
$\vskip-20.00003ptL_{\text{CPR}}=\sum_{\mathbf{y}^{C}\in\mathbb{Y}_{\text{pen}}}\max\Big(0,\;s(\mathbf{y}^{C})-s(\mathbf{y}^{C}_{i,\text{new}})+m\Big),\vskip-20.00003pt$ (7)
where $s(\mathbf{y}^{C})$ denotes the logit assigned to the composition label $\mathbf{y}^{C}$.
Batch-adaptive label-space expansion. Conventional ZS-CAR pipelines optimize classification only over seen compositions $\mathbb{Y}_{\text{seen}}$ (closed-world over compositions), which makes it non-trivial to directly supervise novel verb–object pairs during training. On the other hand, naively applying cross-entropy over the full $\mathbb{Y}^{V}\times\mathbb{Y}^{O}$ space can harm unseen accuracy by treating most unseen compositions as negatives throughout training (see Section 5.4). To provide an effective supervision for novel compositions, we introduce a batch-adaptive label space expansion strategy. For each mini-batch, we construct an expanded label set as $\mathbb{Y}_{\text{exp}}=\mathbb{Y}_{\text{seen}}\cup\mathbb{Y}_{\text{new}}$, where $\mathbb{Y}_{\text{new}}$ denotes the label set for the newly synthesized videos. We compute cross-entropy only over $\mathbb{Y}_{\text{exp}}$. This preserves the optimization stability of closed-world training while progressively injecting supervision for new compositions.
4.2 TORC: Temporal Order Regularization for Composition
To counter object-driven shortcuts in verb learning, we introduce TORC, a temporal order-sensitivity regularizer. While CPR expands supervision over compositions, verb recognition can still collapse to static object cues due to the asymmetric learning difficulty between verbs and objects. As shown in Figure 4 (c), TORC directly regularizes the verb representation to depend on temporal structure: it (i) discourages invariance to temporal reversal by separating forward and reversed features, and (ii) penalizes over-confident verb predictions when temporal order is disrupted by encouraging high-entropy outputs. Together, these constraints reduce reliance on static cues and promote temporally grounded verb representations.
@@ 4.2 TORC: Temporal Order Regularization for Composition
Temporal perturbation for regularization. Given frame-level features $\mathbf{F}=(\mathbf{f}_{1},\dots,\mathbf{f}_{T})$, we form two regularization views: (i) a reversed sequence $\mathbf{F}_{\text{rev}}=(\mathbf{f}_{T},\dots,\mathbf{f}_{1})$, and (ii) a temporally shuffled sequence $\mathbf{F}_{\text{shuffled}}=\pi(\mathbf{F})$, where $\pi$ is a random permutation sampled from $\mathcal{P}$. We then feed $\mathbf{F}_{\text{rev}}$ and $\mathbf{F}_{\text{shuffled}}$ into the verb encoder to obtain perturbed verb features $\mathbf{f}^{V}_{\text{rev}}\in\mathbb{R}^{D}$ and $\mathbf{f}^{V}_{\text{shuffled}}\in\mathbb{R}^{D}$.
@@
TORC loss. We first push the model to distinguish forward and reversed temporal semantics. We minimize the cosine similarity between the original verb $\mathbf{f}^{V}$ and the reversed feature vectors $\mathbf{f}^{V}_{\text{rev}}$: $L_{\cos}=\frac{\mathbf{f^{V}}^{\top}\mathbf{f}^{V}_{\text{rev}}}{\|\mathbf{f}^{V}\|\|\mathbf{f}^{V}_{\text{rev}}\|}$. Temporal reversal often flips action meaning (e.g., opening vs. closing), yet the state-of-the-art model [li2024c2c] yields a high cosine similarity ($0.91$; see Figure 13), indicating weak temporal discrimination.
@@
Next, we regularize the model to avoid confident verb predictions when temporal structure is disrupted. Given the verb text embeddings $\mathbf{E}^{V}=\{\mathbf{e}^{V}_{m}\}_{m=1}^{|\mathbb{Y}^{V}|}$, we define the negative entropy loss $L_{\text{ent}}=\sum_{m=1}^{|\mathbb{Y}^{V}|}p_{m}\log p_{m}$ with the predicted probability $p_{m}$ of the $m$-th verb class based on the cosine similarity between visual and textual features as:
$\vskip-20.00003ptp_{m}$ $=\frac{\exp(\cos(\mathbf{f}_{\text{shuffled}},\mathbf{e}^{V}_{m})/\tau)}{\sum_{n=1}^{|\mathbb{Y}^{V}|}\exp(\cos(\mathbf{f}_{\text{shuffled}},\mathbf{e}^{V}_{n})/\tau)},\vskip-20.00003pt$ (8)
@@
where $\tau$ is a temperature parameter. Maximizing entropy prevents the model from inferring verbs from static spatial cues alone, enforcing reliance on temporal evidence. We define the TORC loss as: $L_{\text{TORC}}=L_{\cos}+L_{\text{ent}}$.
4.3 Training